OpenStudy (anonymous):

What is the radius of a circle in which a 30° arc is 2 pi inches long?

ganeshie8 (ganeshie8):

convert degrees to radians. and use the formula theta = arc-length/radius

OpenStudy (anonymous):

what are radians?

ganeshie8 (ganeshie8):

good question.. you tell me what are degrees ? 360 degrees = 2 pi radians => 30 degrees = pi/6 radians now you can apply theta = l/r formula to get the radius pi/6 = 2 pi/r => r = 12 inches

OpenStudy (anonymous):

wait so would you not use the equation A=n/360 2 pi r? cause thats what it said to use in the book.

OpenStudy (anonymous):

I just could not figure out how to use the that formula to find the radius

ganeshie8 (ganeshie8):

do you understand the formula : theta = (arc length)/radius ?